OpenStudy (anonymous):

On a 10 question multiple-choice test, where each question has four answers, what would be the probability of getting at least one question wrong?

jimthompson5910 (jim_thompson5910):

What's the probability of getting them all correct?

OpenStudy (anonymous):

9.5367x10^-7

jimthompson5910 (jim_thompson5910):

so basically 0.00000095367431

jimthompson5910 (jim_thompson5910):

the probability of getting at least one wrong is one minus this answer because the two events are complementary

jimthompson5910 (jim_thompson5910):

one or the other must happen

jimthompson5910 (jim_thompson5910):

so this means P(getting at least one wrong) = 1-P(getting them all right) P(getting at least one wrong) = 1-0.00000095367431 P(getting at least one wrong) = 0.9999990463257

jimthompson5910 (jim_thompson5910):

so it looks very likely you're going to get at least one wrong

OpenStudy (anonymous):

I thought that that would be right, but I am working the problem for an online class and the computer is saying the answer is incorrect....Maybe its a technicality in how I am entering it.

jimthompson5910 (jim_thompson5910):

do they want it as a decimal? or a fraction?

OpenStudy (anonymous):

all of my other answers have been correct when entered as decimals as long as they are carried out to the fourth decimal place. I have entered 0.9999 and it is saying that is wrong.

jimthompson5910 (jim_thompson5910):

this may sound stupid, but 0.9999990463257 rounded to 4 decimal places is 1.0000

jimthompson5910 (jim_thompson5910):

so that's why I'm thinking they either want more decimal places or they want it as a fraction

jimthompson5910 (jim_thompson5910):

if they want it as a percentage then 0.9999990463257 ---> 0.9999990463257*100 = 99.99990463257% so, 0.9999990463257 = 99.99990463257%

OpenStudy (anonymous):

I tried it as a percentage and that won't work either. I guess I will have to talk with the professor and see what I am doing wrong. Thank you for your help.

jimthompson5910 (jim_thompson5910):

can you post a screenshot of the entire problem so I can have a look?

jimthompson5910 (jim_thompson5910):

there may be some key detail missing

jimthompson5910 (jim_thompson5910):

I'm curious, but have you tried 9999/10000 at all?

OpenStudy (anonymous):

no I haven't tried that

jimthompson5910 (jim_thompson5910):

that's the answer as a fraction, you basically evaluate 1 - (1/(10^4)) to get that

jimthompson5910 (jim_thompson5910):

if neither the decimal or percent forms are working, then it has to be a fraction btw I don't see any piece of info that helps, so maybe there's more context elsewhere, idk?